# Diagnostic-Guided Dynamic Profile Optimization for LLM-based User Simulators in Sequential Recommendation

### v4

@@ Introduction
User simulators (Balog:2024:FnTIR), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (luo2022mindsim). They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive. While early user simulators ranged from rule-based (rohde2018recogym) to deep learning-based approaches (shi2019virtual; zhao2021usersim), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors. To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm. With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (park2023generative) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs. For example, Agent4Rec (zhang2024generative) and RecAgent (wang2025user) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference. By generating the user profile only at initialization, user simulators produce a static representation that cannot adapt to the evolving user interests across multi-round interactions. Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern. Most existing simulators are typically integrated with static recommenders, e.g., MF (koren2009matrix), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
@@
Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation. However, developing this framework faces two primary challenges. (CH1): How to build a reliable mechanism for dynamic profile construction? While existing methods (shu2024rah) like self-reflection (sun2024large) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations. As shown in Fig. 1, directly prompting general-purpose LLMs only achieves an average accuracy of 62.78% on profile defect identification. This highlights the necessity of a more targeted strategy that can accurately diagnose the specific type of defect, such as inaccurate and/or incomplete, and then apply a tailored refinement. To achieve this, we first design a specialized LLM-based diagnostic module through domain-adaptive pre-training and defect-specific fine-tuning, enabling reliable identification of profile defects. As shown in Fig. 1, the module achieves an average accuracy of 92.20%, significantly outperforming the general-purpose LLM. We then design a generalized LLM-based treatment module that uses this precise diagnosis to analyze the defect and generate targeted refinements for the profile. Instead of single-step prompting, our framework iteratively optimizes the profile by traversing the user’s interaction history step-by-step, to dynamically enhance its accuracy and comprehensiveness. (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time? Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (sun2025llm4rsr). This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
Traditional User Simulators for RSs. Early user simulators for RSs progressed from rule-based approaches (rohde2018recogym; ie2019recsim; shi2019pyrecgym) to advanced data-driven techniques (huang2020keeping), including methods based on reinforcement learning (RL) (shi2019virtual; zou2020pseudo), generative adversarial networks (GANs) (chen2019generative; shi2019virtual; zhao2021usersim), and transformers (zhao2023kuaisim; afzali2023usersimcrs). However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
LLM-based User Simulators for RSs. The advent of LLMs has introduced novel paradigms for user simulation in RSs (cai2025agentic; ma2025pub; wei2025mirroring). LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (friedman2023leveraging; zhu2024reliable; zhu2025llm) and generating high-quality interaction data for RL-based algorithms (corecco2024llm; ebrat2024lusifer; zhang2025llm). Current works on user profile construction can be categorized into two types. Some works (zhang2024generative; wang2025user) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences. Others employ self-reflection (zhang2024agentcf; shu2024rah) for iterative refinement. However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
Sequential Recommenders. SRs have evolved from early approaches combining Markov Chains with matrix factorization (rendle2010factorizing), to deep learning models. These include CNNs for local pattern extraction (tang2018personalized), RNNs to capture temporal dependencies (hidasi2015session; zhu2017next; zhu2019query), attention mechanisms to focus on relevant historical items (li2017neural; kang2018self; liu2018stamp; sun2019bert4rec), and GNNs to model high-order relationships within sequential data (wu2022gcrec; yu2024learning; li2024global; liu2024selfgnn). Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
Given a user’s historical interaction sequence $\mathcal{D}_{u}=[i_{1},i_{2},...,i_{T}]$ ordered by timestamp, we split it into two portions: $\mathcal{D}_{u}^{(ini)}=[i_{1},i_{2},...,i_{n}]$, and $\mathcal{D}_{u}^{(opt)}=[i_{n+1},...,i_{T}]$. For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$. It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (zhang2024generative), serving as the foundation user profile. Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@ 9       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (zhang2024generative), which essentially represent a one-time recommendation result split into multiple pages. This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions. To enable realistic multi-round interactions, we propose to incorporate our optimized user simulator DGDPO with sequential recommenders (SRs). Unlike traditional static models, SRs excel at capturing temporal dynamics, which makes them ideal for modeling the bidirectional evolution of user preferences and recommendation strategies over time.
@@
We use three datasets: ML-1M (harper2015movielens), Amazon Books, and Amazon Movies & TV (ni2019justifying). We filter users and items with less than 5 interactions (ML-1M) or less than 20 (Amazon datasets). Interactions are sorted chronologically, with max sequence lengths set to 200 for ML-1M and 50 for Amazon datasets. For each user, the last 10 items are for testing, the 11th most recent item is for validation, and the rest for training (i.e., $\mathcal{D}_{u}$). $\mathcal{D}_{u}$ is further split into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ by an $\alpha:(1-\alpha)$ ratio.
@@
We compare DGDPO with five baselines for user profile construction in two categories. In traditional methods, KuaiSim (zhao2023kuaisim) is a transformer-based model for user profile construction. In LLM-based methods, Agent4Rec (zhang2024generative) utilizes a single-step prompt-based inference. LLM-US (zhang2025llm) generates explanations of user ratings as profile. RAH (shu2024rah) and AgentCF (zhang2024agentcf) employ self-reflection and collaborative reflection, respectively. For multi-round interactions, we integrate the user simulator with four SRs: FPMC (rendle2010factorizing), NARM (li2017neural), SASRec (kang2018self) and LRURec (yue2024linear).
@@
Following state-of-the-arts (shu2024rah; zhang2024generative), we randomly sample 1000 users from each dataset to serve as simulated users. Our diagnostic module is based on lightweight Qwen3-0.6B (yang2025qwen3), while the treatment module employs gpt-4o-mini via its API with a temperature of 0. All LLM-based baselines also utilize GPT-4o-mini as their backbone. For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets. We build upon Agent4Rec (zhang2024generative) as the foundation user simulator and conduct ten-round interactions for each user. For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (sun2022daisyrec). All experiments are conducted on a V100 GPU.
@@
Following (luo2022mindsim; zhang2024generative), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
@@
Fig. 3 depicts the performance of DGDPO across ten interaction rounds on Books dataset under different profile update strategies [^2] Hereafter, we only show the results on Books due to space limitation, and similar trends can be noted on the other two datasets.. The results unveils that dynamic profile update in multi-round interactions can further improve the simualtion accuracy. (1) For ‘w/o Update’, the performance generally exhibits minimal fluctuations across multi-round interactions and is the worst among all strategies. With profile unchanged, ‘w/o Update’ fails to capture the evolution of user preference during multi-round interactions, thereby limiting the simulator’s long-term fidelity. (2)Update w/o GT’ exhibits performance divergence when interacting with different SRs. Specifically, with FPMC, the performance shows a declining trend. In contrast, with advanced deep learning-based SRs, we observe steady performance improvements. This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (hou2024large), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation. (3)Update w/ GT’ performs the best among all methods. This is attributed to an idealized update strategy that always uses the ground truth item to update the user profile, regardless of the simulator’s actions. However, in practical applications, ground truth items are often unavailable, which limits the applicability of this strategy. Therefore, it underscores the importance of developing more advanced user simulators that closely replicate real user behaviors, enabling more effective and reliable updates to the user profile during long-term interactions.

### v5

@@ Introduction
User simulators (Balog and Zhai 2024), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (Luo et al. 2022). They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive. While early user simulators ranged from rule-based (Rohde et al. 2018) to deep learning-based approaches (Shi et al. 2019b; Zhao et al. 2021), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors. To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm. With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (Park et al. 2023) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs. For example, Agent4Rec (Zhang et al. 2024a) and RecAgent (Wang et al. 2025) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference. By generating the user profile only at initialization, user simulators produce a static representation that cannot adapt to the evolving user interests across multi-round interactions. Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern. Most existing simulators are typically integrated with static recommenders, e.g., MF (Koren et al. 2009), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
@@
Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation. However, developing this framework faces two primary challenges. (CH1): How to build a reliable mechanism for dynamic profile construction? While existing methods (Shu et al. 2024) like self-reflection (Sun et al. 2024) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations. As shown in Fig. 1, directly prompting general-purpose LLMs only achieves an average accuracy of 62.78% on profile defect identification. This highlights the necessity of a more targeted strategy that can accurately diagnose the specific type of defect, such as inaccurate and/or incomplete, and then apply a tailored refinement. To achieve this, we first design a specialized LLM-based diagnostic module through domain-adaptive pre-training and defect-specific fine-tuning, enabling reliable identification of profile defects. As shown in Fig. 1, the module achieves an average accuracy of 92.20%, significantly outperforming the general-purpose LLM. We then design a generalized LLM-based treatment module that uses this precise diagnosis to analyze the defect and generate targeted refinements for the profile. Instead of single-step prompting, our framework iteratively optimizes the profile by traversing the user’s interaction history step-by-step, to dynamically enhance its accuracy and comprehensiveness. (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time? Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (Sun et al. 2025). This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
Traditional User Simulators for RSs. Early user simulators for RSs progressed from rule-based approaches (Rohde et al. 2018; Ie et al. 2019; Shi et al. 2019a) to advanced data-driven techniques (Huang et al. 2020), including methods based on reinforcement learning (RL) (Shi et al. 2019b; Zou et al. 2020), generative adversarial networks (GANs) (Chen et al. 2019; Shi et al. 2019b; Zhao et al. 2021), and transformers (Zhao et al. 2023; Afzali et al. 2023). However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
LLM-based User Simulators for RSs. The advent of LLMs has introduced novel paradigms for user simulation in RSs (Cai et al. 2025; Ma et al. 2025; Wei et al. 2025). LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (Friedman et al. 2023; Zhu et al. 2024, 2025) and generating high-quality interaction data for RL-based algorithms (Corecco et al. 2024; Ebrat et al. 2024; Zhang et al. 2025). Current works on user profile construction can be categorized into two types. Some works (Zhang et al. 2024a; Wang et al. 2025) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences. Others employ self-reflection (Zhang et al. 2024b; Shu et al. 2024) for iterative refinement. However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
Sequential Recommenders. SRs have evolved from early approaches combining Markov Chains with matrix factorization (Rendle et al. 2010), to deep learning models. These include CNNs for local pattern extraction (Tang and Wang 2018), RNNs to capture temporal dependencies (Hidasi 2015; Zhu et al. 2017, 2019), attention mechanisms to focus on relevant historical items (Li et al. 2017; Kang and McAuley 2018; Liu et al. 2018; Sun et al. 2019), and GNNs to model high-order relationships within sequential data (Wu et al. 2022; Yu et al. 2024; Li et al. 2024; Liu et al. 2024). Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
Given a user’s historical interaction sequence $\mathcal{D}_{u}=[i_{1},i_{2},...,i_{T}]$ ordered by timestamp, we split it into two portions: $\mathcal{D}_{u}^{(ini)}=[i_{1},i_{2},...,i_{n}]$, and $\mathcal{D}_{u}^{(opt)}=[i_{n+1},...,i_{T}]$. For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$. It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (Zhang et al. 2024a), serving as the foundation user profile. Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@ 9       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (Zhang et al. 2024a), which essentially represent a one-time recommendation result split into multiple pages. This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions. To enable realistic multi-round interactions, we propose to incorporate our optimized user simulator DGDPO with sequential recommenders (SRs). Unlike traditional static models, SRs excel at capturing temporal dynamics, which makes them ideal for modeling the bidirectional evolution of user preferences and recommendation strategies over time.
@@
We use three datasets: ML-1M (Harper and Konstan 2015), Amazon Books, and Amazon Movies & TV (Ni et al. 2019). We filter users and items with less than 5 interactions (ML-1M) or less than 20 (Amazon datasets). Interactions are sorted chronologically, with max sequence lengths set to 200 for ML-1M and 50 for Amazon datasets. For each user, the last 10 items are for testing, the 11th most recent item is for validation, and the rest for training (i.e., $\mathcal{D}_{u}$). $\mathcal{D}_{u}$ is further split into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ by an $\alpha:(1-\alpha)$ ratio.
@@
We compare DGDPO with five baselines for user profile construction in two categories. In traditional methods, KuaiSim (Zhao et al. 2023) is a transformer-based model for user profile construction. In LLM-based methods, Agent4Rec (Zhang et al. 2024a) utilizes a single-step prompt-based inference. LLM-US (Zhang et al. 2025) generates explanations of user ratings as profile. RAH (Shu et al. 2024) and AgentCF (Zhang et al. 2024b) employ self-reflection and collaborative reflection, respectively. For multi-round interactions, we integrate the user simulator with four SRs: FPMC (Rendle et al. 2010), NARM (Li et al. 2017), SASRec (Kang and McAuley 2018) and LRURec (Yue et al. 2024).
@@
Following state-of-the-arts (Shu et al. 2024; Zhang et al. 2024a), we randomly sample 1000 users from each dataset to serve as simulated users. Our diagnostic module is based on lightweight Qwen3-0.6B (Yang et al. 2025), while the treatment module employs gpt-4o-mini via its API with a temperature of 0. All LLM-based baselines also utilize GPT-4o-mini as their backbone. For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets. We build upon Agent4Rec (Zhang et al. 2024a) as the foundation user simulator and conduct ten-round interactions for each user. For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (Sun et al. 2022). All experiments are conducted on a V100 GPU.
@@
Following (Luo et al. 2022; Zhang et al. 2024a), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
@@
Fig. 3 depicts the performance of DGDPO across ten interaction rounds on Books dataset under different profile update strategies [^2] Hereafter, we only show the results on Books due to space limitation, and similar trends can be noted on the other two datasets.. The results unveils that dynamic profile update in multi-round interactions can further improve the simualtion accuracy. (1) For ‘w/o Update’, the performance generally exhibits minimal fluctuations across multi-round interactions and is the worst among all strategies. With profile unchanged, ‘w/o Update’ fails to capture the evolution of user preference during multi-round interactions, thereby limiting the simulator’s long-term fidelity. (2)Update w/o GT’ exhibits performance divergence when interacting with different SRs. Specifically, with FPMC, the performance shows a declining trend. In contrast, with advanced deep learning-based SRs, we observe steady performance improvements. This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (Hou et al. 2024), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation. (3)Update w/ GT’ performs the best among all methods. This is attributed to an idealized update strategy that always uses the ground truth item to update the user profile, regardless of the simulator’s actions. However, in practical applications, ground truth items are often unavailable, which limits the applicability of this strategy. Therefore, it underscores the importance of developing more advanced user simulators that closely replicate real user behaviors, enabling more effective and reliable updates to the user profile during long-term interactions.